# Linying Lv

### v2

@@ 2.2 Information Content
Following Asquith et al. (2005), I use $R^{2}$ as a measure for information content. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
@@
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ is the structured representation of information from analyst report $j$ on stock $i$ released at day $t$. I use the $60\%$ of data as the initial training sample. Therefore, the out-of-sampel test starts in 2015 and ends in 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ as
@@
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
To evaluate the statistical significance of the $R_{\mathrm{oos}}^{2}$, I employ a modified version of the Diebold and Mariano (2002) test for comparing predictive accuracy between two models. Given the strong cross-sectional dependence in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional average of prediction errors from each model rather than individual returns.
@@ 2.3 Information Value
I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal:
@@
where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price of day t-2. $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ denotes the number of analyst reports about stock $i$ on day $t$.
To estimate the price impact $\widehat{\lambda}_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals. The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^2] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression model is
@@
where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda). In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The 3-day interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$. Thus, each regression has $1170(=390*3)$ observations.
Essentially, the measure reflects the ratio of return volatility to price impact. It is derived based on the theoretical framework of Back et al. (2000), which is a continuous-time version of Kyle (1985) model. Due to risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of the asset price. The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance. [^3] Kadan and Manela (forthcoming) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply to the analyst case.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To tackle this issue, I apply the delta method to estimate the mean and variance of information value over sample $s$. The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
2.4 Shapley Value Decomposation
To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology. SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. It was later applied for machine learning model interpretation by Lundberg and Lee (2017) and Chen et al. (2022). In my context, the approach allows me to attribute the overall explanatory power of an analyst report to its constituent topics. Leveraging the additive nature of text embeddings, I design a novel decomposition method to systematically assess the relative importance of various topics in analyst reports.
LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
@@ 2.4 Shapley Value Decomposation
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ is the total number of words in the report. If the tokens can be classified into different topics, the embedding can be fully decomposed into different topic-specific embeddings:
@@
where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
One concern raised by the nature of transformer models is that each token in the input sequence can pay attention to other tokens through a mechanism called “self-attention” or “scaled dot-product attention.” This architecture allows the model to weigh the importance of each token relative to every other token and capture contextual information. For example, the embedding of token $i$ contains information from token $j$, which may pertain to a different topic. To isolate the self-attention across topics, I process the embeddings sentence by sentence to ensure the token embeddings only reflect the context information within the sentence. I then take the token-weighted mean of sentence embeddings as the report representation, which is referred to as sentence-segment embeddings.
@@
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context across sentences.
Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Specifically, the Shapley value can be written as:
@@
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly. [^4] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions.
@@ 3 Data and Empirical Results
This section presents the empirical analysis of analyst report information. I begin by examining the information content and decomposing it across different topics. I then quantify the strategic dollar value of analyst information. The results emphasize analysts’ income statement analyses following earning announcements.
@@ 3.1 Data
Sell-side analyst reports are obtained from the Mergent Investext Database. I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023. The initial reports are in PDF format, and I use Adobe Acrobat to extract plain text from each report. The text is chunked into sentences. To get clean content of firm-relevant discussions, I fine-tune a BERT model to remove the boilerplate segments. [^5] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports. The BERT model fine-tuned for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
@@
The Mergent Investext Database provides information on analyst names and the release date of each report. To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al. (2024). I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database. Since multiple reports could be released between a forecast announcement and its revision, I follow Huang et al. (2014) and define the matching window for each report as 2 days before the announcement date to 2 days after the revision date. [^6] In I/B/E/S, each analyst forecast for a specific firm has two time stamps. The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one. The revision date (REVDATS) is the date the forecast becomes invalid or is revised. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. I also compile firm characteristics from the CRSP and Compustat databases. [^7] See Table A2 and Table A3 for detailed variable definitions and summary statistics. After the matching procedure and excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 stocks from 1,305 analysts across 140 unique brokerage firms.
@@
Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts across years and the Fama-French 12 industries. It also summarizes the average number of pages and tokens per report, which are important indicators of report lengths. Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al. (2023), and the average number of pages per report is consistent with that in Huang et al. (2014). [^8] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline. This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
@@
The stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023. To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
@@
Given my goal to estimate the information value of analyst outputs, I need to control for other information available to investors. In Section 3.2.2, I compare textual information from earnings conference calls and subsequent analyst reports. Earnings conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Available at https://seekingalpha.com/earnings/earnings-call-transcripts.. These transcripts provide comprehensive record of the calls, including executive presentations and manager-analyst interactions like Q&A sessions. I sample earnings conference call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks. This process yielded a final sample of 7,909 firm-quarter-level conference calls.
@@
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ indicates target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $Boldness$ is an indicator variable that takes the value of 1 if revisions are either above both the consensus and previous forecasts or below both. $SR$ represents stock recommendation, following the I/B/E/S rating system, where a rating of 5 denotes a strong sell and a rating of 1 indicates a strong buy. $ERet$ represents the 12-month return forecast, calculated by scaling the 12-month price target by the stock price 50 days before the report release. $Prior\_CAR$ represents the cumulative abnormal return over time period [-10, -2].
Firm-Level Measures: $Size$ is measured by market equity. The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value. Following Huang et al. (2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value. The $Miss$ variable is a binary indicator set to one if $SUE$ is positive. $TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding. The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market fluidity.
Industry-Level Measures: $IndustryRecession$ is an indicator variable equal to one if the firm’s industry (defined by Fama-French 48 classification) experiences a negative monthly return in the bottom quintile.
Macroeconomic Measures: The $TimeTrend$ variable captures the temporal dimension by recording the number of years elapsed since the beginning of the sample period.

### v3

@@ 2.2 Information Content
Following Asquith et al. (2005), I consider $R^{2}$ as a proxy for information content. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations annually using an expanding window:
@@
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ for stock $i$ at day $t$. The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ using:
@@
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the proportion of return variance attributable to analyst information.
To assess the statistical significance of the $R_{\mathrm{oos}}^{2}$, I implement a modified version of the Diebold and Mariano (2002) test for predictive accuracy comparison. Given the strong cross-sectional dependence inherent in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional averages of prediction errors.
@@ 2.3 Information Value
Following Kadan and Manela (2024), I evaluate the value of information using a measure that approximates information value in a strategic setting where multiple informed investors possess imperfect signals:
@@
where $\widehat{\Omega}_{it}$ represents the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price on day t-2. $\widehat{r}_{ijt}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$. Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ represents the number of analyst reports for stock $i$ on day $t$.
To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). The order flow is derived from changes in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals, with trades classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^2] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression specification is:
@@
where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda). In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
The measure reflects the ratio of return volatility to price impact and is grounded in the theoretical framework of Back et al. (2000), which is a continuous-time extension of Kyle (1985) model. Under risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of asset price. The intuition is illustrated with an extended single-auction Kyle model with incomplete information in Online Appendix D.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance. [^3] Kadan and Manela (2024) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply in the analyst context.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To address this challenge, I apply the delta method to estimate the mean and variance of information value over sample $s$. The derivation of the first-order estimator using the delta method is detailed in Online Appendix E.
2.4 Shapley Value Decomposition
To quantify the relative contribution of different topics, I implement the Shapley Additive exPlanations (SHAP) methodology. SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. It was later applied for machine learning model interpretation (Lundberg and Lee, 2017 and Chen et al., 2022). The approach allows for attributing an analyst report’s overall explanatory power to its constituent topics. Leveraging the additive nature of text embeddings, I develop a novel decomposition method to systematically evaluate the relative importance of various topics in analyst reports.
The report-level embeddings are generated as the average of embeddings across all tokens and layers. The report embeddings can be formalized as:
@@ 2.4 Shapley Value Decomposition
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ denotes the total token count of the report. Given the tokens can be classified into distinct topics, the embeddings can be fully decomposed into topic-specific components:
@@
where $P$ denotes the total number of topics, $y_{p}^{emb}$ represents the embedding for topic $p$, $N_{p}$ is the number of tokens classified as topic $p$, and $i_{p}$ indicates the index of tokens within topic $p$.
A consideration in the transformer architectures is the self-attention mechanism, which allows the model to compute relative importance weights between tokens. This architecture enables each token in the input sequence to “talk” with each other. Specifically, the embedding of token $i$ incorporates contextual information from token $j$, even when these tokens belong to distinct topics. To mitigate this cross-topic attention effect, I process embeddings sentence by sentence to ensure the token embeddings only reflect intra-sentence contextual information. I then take the token-weighted mean of sentence embeddings as the report representation, which I designate as sentence-segment embeddings:
@@
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embeddings of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding quantifies the information content of context across sentences.
Subsequently, I compute the Shapley value across topics using sentence-segmented embeddings. The Shapley value is defined as:
@@
where the sum extends over all subsets $S\subseteq P\backslash\{p\}$, including the empty set. $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. The $R_{\text{oos}}^{2}$ is calculated for each unique combination of topic embeddings.  [^4] Shapley value requires a defined input for missing $X$ values. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ across topics, revealing their relative contributions to explaining market reactions.
@@ 3 Data and Empirical Results
In this section, I present the empirical analysis of analyst report information, beginning with an examination of information content and its decomposition across topics. I then quantify the strategic dollar value of analyst information. The results emphasize analysts’ income statement analyses following earning announcements.
@@ 3.1 Data
I refer to the Mergent Investext database for 223,091 sell-side analyst reports for S&P 100 firms spanning the years 2000 to 2023. The reports, initially in PDF format, are converted to plain text using Adobe Acrobat. The extracted text is then segmented into sentences. To ensure a clean dataset focused on firm-relevant discussions, I fine-tune a BERT model to filter out boilerplate content.  [^5] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I construct a training dataset by randomly sampling one report from each of the top 20 brokerage firms by report volume and manually label each sentence. The BERT model fine-tuned for boilerplate classification achieves a 97.31% accuracy rate on test data.
@@
The Mergent Investext Database provides analyst identification and report release dates. To connect Investext with the I/B/E/S database, I first match lead analyst names to their corresponding Analyst IDs (AMASKCD) following Li et al. (2024). I then match individual reports to corresponding EPS forecasts, price targets, and recommendation files in I/B/E/S. Given that multiple reports may be issued between a forecast announcement and its subsequent revision, I adopt Huang et al. (2014)’s approach and define the matching window as spanning from two days before the announcement date to two days after the revision date. [^6] I/B/E/S records two timestamps for each analyst forecast: the announcement date (ANNDATS) when the new forecast is issued and the old one is revised, and the revision date (REVDATS) when the forecast becomes invalid or undergoes revision. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. Additional firm characteristics are sourced from CRSP and Compustat. [^7] Detailed variable definitions and summary statistics are provided in Table A2 and Table A3, respectively. After the matching procedure and excluding reports released on earnings announcement days, the final sample consists of 122,252 reports covering S&P 100 stocks, produced by 1,305 analysts across 140 distinct brokerage firms.
@@
Table 2 provides summary statistics of analyst reports, including the number of reports, unique brokerage firms, and analysts across years and the Fama-French 12 industries. It also reports average page and token counts per report as measures of report length. The temporal distribution of analyst reports aligns with patterns documented in Bonini et al. (2023), while the average page count per report corresponds with findings in Huang et al. (2014). [^8] The data reveals a peak in the number of analyst reports in 2013, followed by a gradual decline. This trend is likely influenced by the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and MiFID II in Europe (Bradshaw et al., 2017).
@@
Stock-level price impacts are estimated using intraday data from the NYSE TAQ database for an out-of-sample period spanning January 2, 2015, to December 28, 2023. To ensure temporal comparability, all prices are adjusted for inflation using the Consumer Price Index (CPI), normalized to December 2020 as the base period.
@@
To isolate the information value of analyst outputs, I control for alternative information sources available to investors. In Section 3.2.2, I conduct a comparative analysis between earnings conference call content and subsequent analyst reports. Conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts.. These transcripts provide detailed documentation of executive presentations and manager-analyst interactions during Q&A sessions. I sample transcripts of S&P 100 constituent firms from 2005 to 2023, matching them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks, yielding a final sample of 7,909 firm-quarter-level conference calls.
@@
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 days pre-release. $Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both. $SR$ denotes stock recommendations following the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell). $ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 days pre-release. $Prior\_CAR$ measures cumulative abnormal return over [-10, -2].
Firm-Level Measures: $Size$ is quantified by market equity. Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value. Following Huang et al. (2018), Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value. $Miss$ is a binary indicator for positive $SUE$. $TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding. Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market fluidity.
Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the firm’s industry (defined by Fama-French 48 classification) experiences negative monthly returns in the bottom quintile.
Macroeconomic Measures: $TimeTrend$ is a temporal variable measuring the number of years since the start of the sample period.